OpenStudy (anonymous):

For every 14 ice-cream cones you buy, you get 2 free. How many free ice-cream cones will you get if you buy 56 cones?

OpenStudy (texaschic101):

14/2 = 56/x (14 to 2 free = 56 to x free) cross multiply 14x = (2)(56) 14x = 112 x = 112/14 x = 8 you will get 8 free when buying 56

OpenStudy (texaschic101):

or you could have just done 56/14 x 2 = 8

OpenStudy (texaschic101):

do you have any questions at all ?

OpenStudy (anonymous):

@texaschic101 this is cool

OpenStudy (texaschic101):

yes...it is :)